Drink more than they should

Gender Yes No Total

Male 151 177 328

Female 92 260 352

Total 243 437 680

What is the relative risk for women thinking they drank more than they should compared to men?
A. 0.41
B. 0.57
C. 1.76
D. 2.41

Please show the steps on how to get the correct answer
Answer #1

Solution :

From given table,

The relative risk for women thinking they drank more than they should compared to men

= (92 / 92 + 260) / 151 / (151 + 177)

= 92 / 352 / 151 / 328

= 30176 / 53152

= 0.57

relative risk = 0.57